OpenStudy (anonymous):

What is the mean of the following data without the outlier? 41, 44, 50, 105

OpenStudy (anonymous):

do you know what an outlier is

OpenStudy (anonymous):

First, identify your outlier. It's the one number that's not close to the other data. Then you remove that number and calculate the mean and avg with the other numbers.

OpenStudy (acxbox22):

in this case 105 is the outlier sice it is not even close to the other numbers 41+44+50=135 135/3=45

OpenStudy (anonymous):

Mean: Average of all the numbers, then divide by how much there is. Median: When you put all the numbers from least to greatest, then you find the middle number. If there are two middle numbers, just average them. Mode: The most frequently shown number. Range: The biggest number minus the smallest number. Outlier: The number that's the farthest away from other numbers. Example: 18,20,17,30. 30 would be the outlier. :) Now to solve your problem, we know that 105 is the outlier. But we need to find mean WITHOUT the outlier. So we count the outlier out. Now, we do 41+44+50=135. There's three numbers, so we do 135/3=45.